Reena is twice old Veena. 3 years ago Reena was 3 time old Veena what age reena
Answer by Cromlix(4381)

Hi there,
Reena is twice old Veena
R = 2V
3 years ago Reena was 3 time old Veena
R - 3 = 3(V - 3)
Replace R with 2V from top equation.
2V - 3 = 3(V - 3)
2V - 3 = 3V - 9
Collect like terms
2V - 3V = - 9 + 3
- V = - 6
V = 6.
Reena is 12 years old
Veena is 6 years old.
Hope this helps :-)